OpenStudy (anonymous):

Suppose a laboratory has a 38 g sample of polonium-210. The half-life of polonium-210 is about 138 days. How many half-lives of polonium-210 occur in 1104 days? How much polonium is in the sample 1104 days later?

OpenStudy (anonymous):

\[m_0\cdot \left(\frac{1}{2}\right)^\frac{t}{t_{1/2}}=m_t \\ 38 \cdot \left(\frac{1}{2}\right)^\frac{1104}{138}=m_{t} \\ m_{t}=38\left(\frac{1}{2}\right)^8=\frac{38}{256}=0.1484 \]

OpenStudy (anonymous):

so 8 ; 0.15 g?

OpenStudy (anonymous):

you round up ?

OpenStudy (anonymous):

I dont think so.

OpenStudy (anonymous):

0.1484375

OpenStudy (anonymous):

answers are: 9 ; 0.07 g 8 ; 2,2622 g 8 ; 4.75 g 8 ; 0.15 g

OpenStudy (anonymous):

8 ; 0.15 g rounding up then